# Exploring Self-Supervised Skeleton-Based Human Action Recognition under Occlusions

### v2

@@ III-B Evaluation of KNN Imputation
TABLE I: Linear evaluation results on NTU-60 with synthesized realistic occlusion, randomly imputed values, and imputed values by our proposed method. “$\Delta$” represents the difference compared to the non-imputed NTU-60. J and M represent the joint stream and the motion stream.
Method Stream Occluded (%) Randomly imputed (%) Our imputed (%)
@@
TABLE II: Linear evaluation results on NTU-120 dataset.
Method Stream Occluded (%) Randomly imputed (%) Our imputed (%)
@@
TABLE III: Finetune and Semi-supervised results on the imputed NTU-60/120 datasets.
@@
III-C Evaluation of OPSTL framework
We conduct a series of comparison experiments to evaluate OPSTL on this new challenge proposed by this work. As shown in Tables III, and III, for all the non-imputed, randomly imputed, and KNN-imputed NTU-60/120, OPSTL outperforms the current state-of-the-art method PSTL on linear evaluation. OPSTL is improved by $1.59\%$ and $1.95\%$ on xsub and xview of NTU-60 with realistic occlusion, respectively. In addition, OPSTL achieves a performance gain of $1.47\%$ and $2.28\%$ on xsub and xset of NTU-120 with realistic occlusion. OPSTL achieves not only improvements on the non-imputed dataset but also on the imputed NTU-60/120 dataset.
III-D Ablation Study
1) To validate the effectiveness of the imputation method, we perform random imputation on NTU-60/120 with realistic occlusion. From Table I and Table II, it is obvious that our proposed imputation method outperforms random imputation on linear evaluation. The performances of SkletonCLR, 2s-CrosSCLR, and AimCLR all deteriorate under random imputation. The most severe decline is observed in AimCLR, where the accuracy drops to nearly $1\%$. This clearly illustrates that random completion undermines the whole skeleton information, which is crucial for action recognition methods that leverage whole skeleton data.
On the contrary, in the case of methods that utilize partial skeleton sequences for representation learning, such as PSTL, there is still a slight improvement in accuracy despite the use of random imputation. However, for our proposed ASM, the improvement is minimal or even negative. For instance, on xset of randomly imputed NTU-120, OPSTL experiences a reduction of $0.28\%$ in accuracy compared to non-imputed data. Yet, it still outperforms the state-of-the-art PSTL. This observation underscores the fact ASM effectively utilizes high-quality data for representation learning.
TABLE IV: Stepwise ablation results on realistic occluded and imputed NTU-120. Method1+2 denotes two stages during the pre-training. S1 is first stage and S2 is second stage. All experiments are on the joint stream.
@@ III-D Ablation Study
2) To better illustrate the effectiveness of ASM, we conducted a stepwise ablation study on the NTU-120 dataset. As shown in Table IV, during the first stage of pre-training, we observed that using ASM yields accuracy improvements of $1.47\%$ for xsub and $2.28\%$ for xset over CSM. Building upon the ASM-based first stage, in the second stage, both CSM and ASM are employed. The results indicate in the second stage, continuing to use ASM yields gains of $0.59\%$ for xsub and $0.73\%$ for xset compared to using CSM.
3) To demonstrate the effectiveness of our imputation method on random occlusions, we validate it using OPSTL on the NTU-60/120 datasets. As shown in Table V, after the imputation, OPSTL exhibits significant improvements across various splits of the NTU-60/120 datasets.
TABLE V: Linear evaluation results of OPSTL on non-imputed and imputed NTU-60/120 with random occlusion. All experiments are on the joint stream.
@@
IV Conclusion
In this paper, we propose effective solutions for the challenges of self-supervised skeleton-based action recognition in occluded environments. First, we construct a large-scale occluded self-supervised skeleton-based human action recognition benchmark considering well-established approaches and propose the Adaptive Spatial Masking (ASM) data augmentation mechanism to formulate OPSTL framework. We introduce a skeleton data imputation method using KMeans and KNN, reducing computation overhead for occluded skeleton completion, named as KNN imputation. Experimental results validate the efficacy of our approach across various self-supervised approaches, empowering robots to perform robust action recognition in real-world occluded scenarios.

### v3

@@ IV-A Datasets
TABLE II: Comparison among no imputation, random imputation, and our proposed cluster agnostic KNN imputation on various of approaches. Linear evaluation results on NTU-60 with realistic synthesized occlusion are reported.“$\Delta$” represents the difference compared to the non-imputed NTU-60. J and M represent the joint stream and the motion stream. Note that, our contributed method incorporates both the OPSTL and the new imputation method.
Method Stream No imputation (%) Random imputation (%) Our imputed (%)
@@
TABLE III: Comparison among no imputation, random imputation, and our proposed cluster agnostic KNN imputation on various of approaches. Linear evaluation results on NTU-120 with synthesized realistic occlusion are reported. “$\Delta$” represents the difference compared to the non-imputed NTU-120. J and M represent the joint stream and the motion stream.
Method Stream No imputation (%) Random imputation (%) Our imputation (%)
@@
IV-B Protocols
Linear Evaluation Protocol. To elaborate, we train a supervised linear classifier consisting of a fully connected layer followed by a softmax layer while keeping the encoder fixed.
Semi-supervised Evaluation Protocol. We initially pre-train the encoder using the entire imputed dataset and subsequently fine-tune the complete model using only $1\%$ or $10\%$ randomly chosen labeled data.
Finetune Protocol. We attach a linear classifier to the trained encoder and finetune the entire network on the imputed data.
IV-C Implementation Details
In our experiments, all pre-trained models are based on ST-GCN [27] with $16$ hidden channels as the backbone. The preprocessing steps closely follow those used in CrosSCLR and AimCLR. This involves removing invalid frames from skeleton sequences, resizing sequences to $50$ frames using linear interpolation, and transforming coordinates into relative coordinates. We also compute the distribution of missing joints ($\mathbf{B}$) in the data.
For training, we employ the Adam optimizer [42] and use the CosineAnnealing scheduler with a total of $150$ epochs for both representation learning and downstream tasks. Our batch size is set to $128$. The learning rate is set to $35e{-}$ 3.
Data Augmentation. Data augmentation is performed to diversify skeleton sequences before feature extraction during model training. Each model uses its specific set of data augmentation methods. For instance, SkeletonCLR and CrosSCLR utilize one spatial augmentation (Shear) and one temporal augmentation (Crop). AimCLR employs four spatial augmentations (Shear, Spatial Flip, Rotate, Axis Mask) and two temporal augmentations (Crop, Temporal Flip), along with two spatiotemporal augmentations (Gaussian Noise and Gaussian Blur). PSTL uses three spatial augmentations (Shear, Rotate, Spatial Flip) and one temporal augmentation (Crop).
Self-supervised Pre-training. To ensure a direct comparison with PSTL, we use the same set of parameters. As shown in Fig. 2, the incomplete skeleton sequence $S$ through transformation $\tau$ has three different views $x$, $x^{\prime}$, $\hat{x}$. $x^{\prime}$ and $\hat{x}$ pass through ASM and MATM to create partial skeleton sequence $x^{\prime}_{a}$, $\hat{x}_{m}$. ST-GCN is used as the shared backbone $f$ to extract $256$-dimensional features $h$, $h^{\prime}_{a}$, $\hat{h}_{m}$, which are then projected to $6144$-dimensional embeddings $z$, $z^{\prime}_{a}$, $\hat{z}_{m}$ through shared projector $g$. To capture the relationship between masked joints and unmasked ones, we compute the cross-correlation matrix between embeddings $z$ and $\hat{z}_{m}$, as well as $z$ and $z^{\prime}_{a}$. $L1$ and $L2$ are calculated using the two cross-correlation matrices. The loss parameter $\lambda$ is set to $2e^{-4}$, and a warm-up strategy of $10$ training epochs is applied. Weight decay is set to $1e^{-5}$. For ASM, $9$ joints are masked, and for MATM, $10$ frames are masked.
Imputation. We proposed an imputation method to deal with occlusion. For clustering during imputation, we employ KMeans with $60$ clusters for NTU-60 and $120$ clusters for NTU-120 with realistic occlusions. We use KNN with a value of $k$ set to $5$ to search for neighboring samples in the imputation process. The chosen cluster number corresponds to the number of the total class.
IV-D Benchmark Analysis
We propose Imputation Occluded Skeleton from Partial Spatial-Temporal Learning (IosPSTL) to address the occlusion challenge in self-supervised action recognition. As shown in Table I, our approach achieves significant improvements on occluded datasets compared with the selected baselines, with approximately $8\%$ and $6\%$ gains on the occluded NTU-60 and NTU-120, respectively, compared to other methods. Then we would like to examine the effectiveness of the proposed cluster-agnostic KNN imputer on other self-supervised skeleton-based human action recognition baselines. We compare all the baselines together with our proposed OPSTL w/o imputer, w/ random imputer, and w/ cluster agnostic KNN imputer on NTU-60/120 with realistic occlusion. As shown in Tables IIIII, and IV, with the help of the proposed cluster agnostic KNN imputer, almost all three downstream task performances of all the investigated methods have shown obvious improvements on both of the NTU-60 and NTU-120 datasets.
The performance gains brought by our proposed imputer are demonstrated on the right hand side of Table II and marked by $\Delta$.
OPSTL achieves a notable enhancement of $6\%$ and $5.84\%$ in xsub and xview of imputed NTU60, respectively. AimCLR demonstrates improvements of $7.92\%$ and $3.90\%$ in xsub and xset of imputed NTU120, respectively. The enhancement in accuracy after imputation is indicated by red upward arrows.
TABLE IV: Finetune and Semi-supervised results on the imputed NTU-60/120 with synthesized realistic occlusion.“$\Delta$” represents the difference compared to the non-imputed NTU-60/120 with synthesized realistic occlusion. J and M represent the joint stream and the motion stream.
@@ IV-D Benchmark Analysis
Additionally, we conduct a series of comparison experiments to evaluate OPSTL. As shown in Tables IIIII, and IV, for all the non-imputed, randomly imputed, and imputed NTU-60/120, OPSTL outperforms the most performing baseline, i.e., PSTL on linear evaluation. The proposed ASM mechanism delivers performance improvements of $1.59\%$ and $1.95\%$ on xsub and xview of NTU-60 with realistic occlusion, respectively. In addition, OPSTL achieves a performance gain of $1.47\%$ and $2.28\%$ on xsub and xset of NTU-120 with realistic occlusion. Not only on the non-imputed dataset but also on xsub and xset of imputed NTU-120 achieves $2.24\%$ and $0.31\%$ improvements.
IV-E Ablation Study
We conduct ablation experiments to demonstrate the effectiveness of the proposed adaptive Spatial Masking (ASM) and the cluster agnostic KNN imputer method.
1) To validate the effectiveness of the imputation method, we perform random imputation on NTU-60/120 with realistic occlusion. As shown in the Table II, the cluster-agnostic KNN imputer achieves $4.50\%$ and $1.47\%$ performance improvements (shown on the right-hand side) on the baseline AimCLR when comparing with the AimCLR without imputation. However, random imputation can not achieve desired performance improvements, resulting in $52.54\%$ and $3.30$ performance decays. This observation indicates that the cluster agnostic KNN imputer can achieve better performance than random skeleton imputation.
The experiments also revealed that random imputation is beneficial for models trained using a partial skeleton sequence but detrimental for models trained using a complete skeleton sequence, leading to performance degradation. This is because randomly imputed skeleton data provides compensatory information that is useful for model learning from partial skeleton data. Therefore, it is necessary to conduct further research on random imputation. From Table II and Table III, it is obvious that our proposed cluster agnostic KNN imputer outperforms random imputation on linear evaluation. The performances of SkeletonCLR, 2s-CrosSCLR, and AimCLR all deteriorate under random imputation.
TABLE V: Stepwise ablation results on realistic occluded and imputed NTU-120. Method1+2 denotes two stages during the pre-training. S1 is first stage and S2 is second stage. All experiments are on the joint stream.
@@ IV-E Ablation Study
2) To better illustrate the effectiveness of ASM, we conducted a stepwise ablation study on the NTU-120 dataset. As shown in Table V, during the first stage of pre-training, we observed that using ASM yields accuracy improvements of $1.47\%$ for xsub and $2.28\%$ for xset over CSM. Building upon the ASM-based first stage, in the second stage, both CSM and ASM are employed. The results indicate in the second stage, continuing to use ASM yields gains of $0.59\%$ for xsub and $0.73\%$ for xset compared to using CSM.
3) To demonstrate the effectiveness of our imputation method on random occlusions, we validate it using OPSTL on the NTU-60/120 datasets. As shown in Table VI, after the imputation, OPSTL exhibits significant improvements across various splits of the NTU-60/120 datasets.
TABLE VI: Linear evaluation results of OPSTL on non-imputed and imputed NTU-60/120 with random occlusion. All experiments are on the joint stream.
@@
V Conclusion
In this paper, we propose effective solutions for the challenges of self-supervised skeleton-based action recognition in occluded environments. First, we construct a large-scale occluded self-supervised skeleton-based human action recognition benchmark considering well-established approaches. Due to the limited performances of the leveraged baselines, we propose a new approach, named IosPSTL, which delivered state-of-the-art performances on the NTU-60/120 datasets. As part of the IosPSTL, cluster agnostic KNN imputer method using KMeans and KNN, reducing computation overhead for occluded skeleton completion is proposed. Apart from the data-driven approach, OPSTL with ASM is proposed to improve the robustness of feature learning during the contrastive training. Experimental results validate the efficacy of our approach across various self-supervised skeleton-based human action recognition approaches, empowering robots to perform robust action recognition in real-world occluded scenarios.